If someone could just give me an example as to how I can solve the below? Am I suppose to enter 1,2,4,8 and 16 for x on each one? Just a detailed example of one or a few of these would help me so much. Just something to get me started.


"Evaluate each of the functions below at x = 1, 2, 4, 8, and 16. Plot the graph of each function. Classify each as linear, quadratic, polynomial, exponential, or logarithmic, and explain the reasons for your classifications. Compare how quickly each function increases, based on the evaluations and graphs, and rank the functions from fastest to slowest growing."
f(x) = x^3 - 3x^2 - 2^x + 1
f(x) = e^x
f(x) = 3x - 2
f(x) = log x
f(x) = x^2 - 5x + 6

Answer by stanbon(57337)

Yes, substitute the number, like x=2, in place of the x, as follows:
-----------
f(x) = x^3 - 3x^2 - 2^x + 1
f(2)=2^3 - 3*2^2 -2^2 +1
=8 - 12 - 4 +1
=-7
The function is the sum of a and an exponential.
----------
f(x) = e^x
f(2) = e^2 = 7.389056...
The function is exponential because the "x" is in the exponent.
----------------------
f(x) = 3x - 2
f(2) = 3*2 -2 = 4
The function is linear ; it is a polynominal
------------------
f(x) = log x
f(2) = log 2 = 0.030102999...
The function is logrithmic
------------------
f(x) = x^2 - 5x + 6
f(2) = 2^2 -5*2 +6 = 4 -10 +6= 0
The function is quadratic; it is a polynomial
-----------
Cheers,
Stan H.